OpenStudy (anonymous):

A rectangular is five times as long as it is wide. If the area is 80ft^2, find the length and the width of the table.

OpenStudy (anonymous):

x = width 5x = length x * 5x = 80 5x^2 = 80 x^2 = 16 x = -4, 4 since measurement has to be positive it's 4 width is 4 length is 20